OpenStudy (anonymous):

MEDALS FOR BEST ANSWERS Solve for x: 3 − (2x − 5) < −4(x + 2) Solve for x: −2x − 4 > 8

OpenStudy (anonymous):

you can substitute the <, and > signs for an = sign

OpenStudy (anonymous):

so 3-(2x-5) = -4(x+2)

OpenStudy (anonymous):

then you will distribute on both sides

OpenStudy (anonymous):

what do you come up with

OpenStudy (anonymous):

3-2x+5 = ??? (your turn)

OpenStudy (anonymous):

6?

OpenStudy (anonymous):

3-2x+5=-4x-8

OpenStudy (anonymous):

then combine like terms 2x+8=-8

OpenStudy (anonymous):

x < −8 x > −8?

OpenStudy (anonymous):

2x=0 x=0 x<0

OpenStudy (anonymous):

Solve for x: −2x − 4 > 8

OpenStudy (anonymous):

-2x-4>8 -2x-4=8 -2x=12 -2x>12 when dividing by a negative you will flip the sign x<-6

OpenStudy (anonymous):

do u understand what i did?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

:)